# Theoretical study on the crystal structure of a bilayer nickel-oxychloride Sr3Ni2O5Cl2 and analysis on the occurrence of possible unconventional superconductivity

### v1

@@ III.4 Formation energy (enthalpy)
To examine the possibility for the synthesis of Sr3Ni2O5Cl2, we calculated formation energies for some possible chemical reactions as presented in Table 2. Here, the formation energy is defined as the total energy of the right-hand side of the chemical equation relative to that for the left-hand side. We found that Sr2NiO2Cl2 + SrNiO3 and SrNiO3 + SrNiO2 + SrCl2 have similar total energies to that of the target compound Sr3Ni2O5Cl2. This result suggests that, even by starting from the relatively unstable compounds including SrO2, which was used in synthesis of similar oxychlorides, Sr2NiO2Cl2 [157], Sr2CoO3Cl, and Sr3Co2O5Cl2 [139], a synthesis can be trapped in the middle of the reaction.
However, this problem might be resolved by the high-pressure synthesis, as adopted in Ref. 157 for Sr2NiO2Cl2. Figure 5 presents that the formation enthalpy with respect to the applied pressure for chemical reactions (c)(d) that are defined in Table 2. Here, the enthalpy is defined as the sum of the total energy and the $pV$ term, where $p$ and $V$ are the pressure and the volume, respectively. We found that the formation enthalpy goes to negative even for reaction (d).
@@ Acknowledgements.
Figure 6, presents the calculated phonon dispersion for La3Ni2O7 under 0 and 20 GPa. We confirmed that the crystal structure with the $I4/mmm$ space group is unstable without applying the pressure but is stable by applying a sufficient pressure as reported in several studies [38, 55, 54].

### v2

@@ III.4 Formation energy (enthalpy)
To examine the possibility for the synthesis of Sr3Ni2O5Cl2, we calculated formation energies for some possible chemical reactions as presented in Table 2. Here, the formation energy is defined as the total energy of the right-hand side of the chemical equation relative to that for the left-hand side. Huge (negative) formation energies for reactions (a)(b) mean that the left-hand sides of (a)(b) are very unstable and unlikely the final products after the chemical reaction, while they are perhaps good starting points to promote the chemical reaction. On the other hand, we found that Sr2NiO2Cl2 + SrNiO3 and SrNiO3 + SrNiO2 + SrCl2 have similar total energies to that of the target compound Sr3Ni2O5Cl2. This result suggests that, even by starting from the relatively unstable compounds including SrO2, which was used in synthesis of similar oxychlorides, Sr2NiO2Cl2 [100], Sr2CoO3Cl, and Sr3Co2O5Cl2 [80], a synthesis can be trapped in the middle of the reaction.
However, this problem might be resolved by the high-pressure synthesis, as adopted in Ref. 100 for Sr2NiO2Cl2. Figure 5 presents that the formation enthalpy with respect to the applied pressure for chemical reactions (c)(d) that are defined in Table 2. Here, the enthalpy is defined as the sum of the total energy and the $pV$ term, where $p$ and $V$ are the pressure and the volume, respectively. We found that the formation enthalpy goes to negative even for reaction (d).
@@ Acknowledgements.
Figure 6 presents the calculated phonon dispersion for La3Ni2O7 under 0 and 20 GPa. We confirmed that the crystal structure with the $I4/mmm$ space group is unstable without applying the pressure but is stable by applying a sufficient pressure as reported in several studies [32, 42, 41].
@@
Appendix D: cRPA interaction parameters for Sr3Ni2O5Cl2
We evaluated the onsite interaction parameters for the ten-orbital model of Sr3Ni2O5Cl2 using the cRPA [105] with the projector method [106]. For this purpose, we used the plane-wave cutoff energy of 450 eV, 200 bands including occupied and unoccupied states, and an $8\times 8\times 8$ ${\bm{k}}$-mesh. Obtained interaction parameters are shown in Table 3.
Table 3: Interaction parameters of onsite Coulomb repulsion $U,U^{\prime}$ and Hund’s coupling (pair-hopping) $J$ ($J^{\prime}$) evaluated with cRPA. The orbital indices $l,m=1,2,3,4,5$ indicate the $d_{x^{2}-y^{2}},d_{3z^{2}-r^{2}},d_{xy},d_{yz},d_{zx}$ orbitals, respectively. For simplicity, we abbreviate four-orbital-index-representation of partially screened interaction integrals $V$, namely, $V_{llmm}\rightarrow U^{\prime}_{lm}$ and $V_{lmlm}(=V_{lmml})\rightarrow J_{lm}$. Note that $U^{\prime}_{lm}=U^{\prime}_{ml}$, $J_{lm}=J_{ml}$, and $J_{lm}=J^{\prime}_{lm}$. All in eV.
$l,m$ $U,U^{\prime}$ $J,J^{\prime}$
$1,1$ 3.936   —
$1,2$ 2.866 0.563
$1,3$ 3.151 0.249
$1,4$ 2.990 0.473
$1,5$ 2.990 0.473
$2,2$ 4.080   —
$2,3$ 2.683 0.460
$2,4$ 3.365 0.347
$2,5$ 3.365 0.347
$3,3$ 3.393   —
$3,4$ 2.802 0.420
$3,5$ 2.802 0.420
$4,4$ 4.088   —
$4,5$ 3.040 0.472
$5,5$ 4.088   —